David Z. Maze

103 Willow Ave. #2 Somerville, MA 02144-2527 (617) 625-2181

                                         dmaze@mit.edu

Objective

To find a permanent position in the field of computer science, with particular interest in research and development of compilers.

Experience

International Business Machines (DataPower Technology) . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Cambridge, MA Software Engineer February 2004-present Worked on XML parser and compiler and runtime for XSLT, a functional XML-to-XML transformation language. Designed and implemented a W3C XML Schema validator within this framework. Built parser and schema validator in conjunction with hardware group to run on locally developed PCI card.

Massachusetts Institute of Technology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Cambridge, MA Research Scientist September 2001-February 2004 Worked principally on StreamIt, a language and compiler for streaming applications, including work on a uniprocessor runtime library, a compiler front-end, and general work within the compiler. Other tasks included general system administration, installation and maintenance of a bug-tracking system, and design and implementation of an online paper database.

Akamai Technologies . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Cambridge, MA Software Intern October 1999-April 2001 Enhanced build infrastructure for source code. Developed Web-based administration tools for Perforce source control management system. Assisted in deployment of automated testing infrastructure.

Altera Corporation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . San Jose, CA Software Intern May 1999-August 1999 Added an optimization stage to a hardware description language compiler for programmable logic devices.

Education

Massachusetts Institute of Technology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Cambridge, MA Received Master of Engineering in Electrical Engineering and Computer Science September, 2001 Developed a compilation and simulation infrastructure to compare the performance of VLIW and SIMD architectures across a common application set.

Massachusetts Institute of Technology . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . Cambridge, MA Received Bachelor of Science in Electrical Engineering and Computer Science June, 2000

Proficiencies

Computer skills include eight years of experience with UNIX, and familiarity with MacOS, Windows, Linux, Solaris, IRIX. Extensive knowledge of C, Java, Perl, Python, XSLT, and standard UNIX and XML tools. Programming experience also includes C++, PostScript, Lisp, Haskell, and SQL.

Other Activities

Debian GNU/Linux . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . February 2001-October 2003 Package maintainer for xcircuit schematic editor and lm-sensors hardware status monitor. Tasks include tracking and packaging original sources, fixing bugs, and resolving conflicts with other packages.

Student Information Processing Board . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . September 1999-February 2004 SIPB provides software, support, and computer services to users of MIT’s Athena Computing Environment. Member-at-large of the SIPB Executive Committee from February 2000 to September 2000. Secretary from September 2000 to December 2000.